# Universal Scaling Laws of Absorbing Phase Transitions in Artificial Deep Neural Networks

### v2

@@ IV.3 Scaling results for the finite networks and different architectures
Figure 3: Finite-size scaling of the order-to-chaos transition in the multilayer perceptrons (1). (a) The order parameter $\rho^{(l)}$ (see Eq. (39)) at a critical point ($(\sigma_{w;c},\sigma_{b})\sim(1.39558,0.3)$ with $\tanh$ activation; $\kappa\sim 0.233498$) for various widths $n$ (ranging from 50 (purple) to 400 (orange)), empirically averaged over $10^{4}$ independent runs. Two orthogonal inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ (with $\|\boldsymbol{x}_{1}\|_{2}=\|\boldsymbol{x}_{2}\|_{2}=1$) of size $n_{\mathrm{in}}=10$ were given. The inset shows the same data rescaled according to the univesal scaling ansatz (40). The black dashed curve indicates the solution (42) of the phenomenological description (41) with $(\kappa,\mu)\sim(0.233498,0.6601)$, where $\mu$ was chosen by fitting the solution (42) to the empirical result for $n=400$. (b) The nonuniversal metric factor $\mu$ as a function of $\kappa$ in the case of $\tanh$ activation, where $\mu$ for each $\kappa$ was estimated from the same fitting as the inset of (a) using the empirical $\rho^{(l)}$ for $n=200$.
The empirical finding above can be heuristically understood by considering a finite-width correction to the mean-field theory. In the case of the finite-width networks, the fourth-order (and other even-order) cumulants come into play, while the third-order (and other odd-order, except the first) ones vanish  [^3] This peculiarity explains why the finite size scaling in the multilayer perceptrons is different from that in the contact process [83] on a complete graph, where one finds the same $\beta$ and $\nu_{\|}$ (Eq. (21)) but the exponent for finite-size scaling (40) is replaced with $-1/2$. If the third-order cumulant remained non-zero, the leading order for the correction would be an order of $n^{-\frac{1}{2}}$.. In the spirit of the asymptotic expansion of the probability distribution [43], this observation indicates that the leading correction to the Gaussian process is an order of $n^{-1}$, the reciprocal of the width. Thus, together with an obvious fact that $\rho=0$ is an absorbing state also for the finite networks, we are led to the following modified phenomenological description:
@@
While it is nowadays well established that the depth-to-width ratio $L/n$ is a key quantity for describing the multilayer perceptrons with finite $n$ [23, 44], the metric factor $\mu$ enriches this insight by providing the means to quantitatively characterize the sensitivity of the network to the width. Specifically, the width $n$ of the network should satisfy $n\gtrsim\mu L$ so that the signal propagation dynamics therein is reasonably well approximated by the infinitely wide limit. This introduces another design consideration for the neural networks. Recalling the relevance of $\kappa L$ for the training dynamics of the networks, one would be tempted to use a critical network with larger $\kappa$ to obtain good generalization with smaller $L$. However, Fig. 3(b) empirically suggests that larger $\kappa$ comes with a cost of larger $\mu$, which imposes an extra computational burden for larger $n$. Hence, if one chooses to operate the network near the infinitely wide limit, one needs to make a trade-off between these two factors for cost-effective training. Theoretically, a more precise formulation of this idea might be achieved by studying the finite-width corrections [45] to the training dynamics, which is beyond the scope of the present work.
@@
although some deviation is well expected due to several reasons, such as the finite-size effect, deviation of the noise from the asymptotic behavior (which is particularly true if $c$ is small), and the saturation to non-zero order parameter slightly below the critical point, just like DP with a small external field. We also checked that the essentially same scenario holds true for the two-dimensional convolutional networks (Fig. 4(b)), where the critical exponents $\beta,\nu_{\|}$ are replaced [53, 54] with
@@
These empirical observations suggest that the signal propagation dynamics in the convolutional networks is highly nontrivial, especially given the notorious difficulty of exactly solving DP [55]. Yet, thanks to the universality of the scaling laws of absorbing phase transitions, semi-quantitative predictions can be gained via simple phenomenological considerations. In particular, the most informative combination of the width $n$ and the depth $L$ for describing the behavior of the critically initialized deep convolutional networks may be changed from $L/n$ to $L/n^{\nu_{\|}/\nu_{\perp}}$ using the corresponding critical exponents
@@ V Discussion
Let us emphasize that successful deep learning can only be achieved via a complicated interplay among various setups, even in one of the simplest cases where NTK describes the training dynamics reasonably well. In addition to the relevance of initialization at criticality [18, 21] and of proper scaling of a learning rate with respect to depth [19], the present work demonstrates the necessity of a more specific, depth-dependent choice of hyperparameters. Furthermore, in the case of large but finite width, one should also strike a balance between width and depth for efficiency. The fact that all these setups need to be considered simultaneously highlights the major challenge in deep learning, which necessitates extensive study on hyperparameter optimization [58]. To put it the other way around, considerable theoretical insight into the mechanism behind the recent success of deep learning may be obtained by studying the neural networks near the realm of NTK, in contrary to a common belief [59, 60].
@@
We foresee some interesting directions for future work. Apart from the ones mentioned in the previous Section  [^4] Namely, (a) derivation of more precise design principles of the neural networks from the universal scaling laws, possibly with prior knowledge about a dataset at hand (Sect. IV.2) and (b) uncovering the details of the trade-off between the width and the depth (Sect. IV.3)., one of the most natural directions is to extend the present framework to more modern architectures. In particular, the skip connections employed in the residual neural networks (ResNet [69]) may be seen as a stimulus kicking the system out of an absorbing state: even if the signals are collapsed during the propagation within a residual block, the skip connection breaks the order before entering the new block. Given that a combination of external drive and dissipation into absorbing states has been conjectured to be a key ingredient of self-organized criticality in physical systems [3, 7], the skip connections may have their unique benefits in deep learning as well. Another, albeit less straightforward, direction is to provide thermodynamic foundations of the speed-accuracy trade-off in deep learning. The deep neural networks in the ordered phase sacrifice speed for accuracy, and vice versa in the chaotic phase. Since the speed-accuracy trade-off has been extensively studied in the context of living systems [70, 71, 72, 73], thermodynamical insights developed therein are likely to be helpful (and indeed, very recently, the trade-off in the diffusion models has been studied from a thermodynamic viewpoint [74]), although pursuing this direction would call for an improved understanding of the thermodynamics of absorbing phase transitions [75, 76]. We believe that further investigation into a parallel between intelligence in living systems and that in artificial neural networks will lead us to a lot of exciting developments, beneficial for both physics and machine learning communities.
@@ Appendix B Scale-invariant activation functions
$(\sigma_{w},\sigma_{b})=\left(\sqrt{\frac{2}{1+a^{2}}},0\right).$ (68)
Note that a special case of $a=0$ is nothing but the well-known He initialization [77] for ReLU activation. With this initialization scheme, we have a well-defined iterative $\mathcal{C}$-map (12) and hence we can study the order parameter $\rho^{(l)}$ defined in Eq. (20); Cho and Saul [78] provide technical details on how to analytically deal with the integration appearing in Eq. (12) in the case of ReLU activation. Notice also that $q^{*}$ vanishes for $\sigma_{w}<\sigma_{w;c}$ if $\sigma_{b}=0$. In other words, the neurons within sufficiently deep hidden layers die out [79], which is reminiscent of an absorbing phase transition discussed in the main text.
Figure 5: Universal scaling in the infinitely-wide multilayer perceptrons (1) with scale-invariant activation functions (67). (a) The order parameter $\rho^{(l)}$ at the edge of chaos (68) for various cosine distances $\rho^{(0)}$ (ranging from $10^{-3}$ to $10^{-1}$; see Eq. (26)) of the inputs and leak parameters $a$, calculated from the mean-field theory (6), (7) and then rescaled according to the scaling ansatz (69). Special cases of $a=0$ and $a=1$ correspond to ReLU and linear functions, respectively. The dashed lines are guides-to-eye for the asymptotic behavior of the scaling function $g$. The inset shows the nonuniversal metric factor (72) as a function of $a$. (b) The NTK $\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})$ (see Eq. (32)) for various network depths, rescaled according to the universal scaling ansatz (70). The two horizontal lines are guides-to-eye for the asymptotic behavior as $L\rightarrow\infty$ [21]: the upper one for $\boldsymbol{x}_{1}=\boldsymbol{x}_{2}$ ($\rho^{(0)}=0$), the lower one for $\boldsymbol{x}_{1}\neq\boldsymbol{x}_{2}$ ($\rho^{(0)}>0$).
A natural question is whether one can apply the universal scaling of absorbing phase transitions in the present case, which we will address below. As visualized in Fig. 5(a), we find that the order parameter $\rho^{(l)}$ at the edge of chaos (68) follows the universal scaling ansatz
$\rho^{(l)}\simeq(\kappa l)^{-2}g(\rho^{(0)}(\kappa l)^{2}),$ (69)
where we dropped the metric factor $\omega$ associated with an initial condition because $\omega=1$ in this case. Similarly, the universal scaling for NTK holds for small $\rho^{(0)}$ or large $L$ (Fig. 5(b)):
@@
The difference in the scaling exponent compared to the result (27) in the main text stems from the second-dominant term in the iterative $\mathcal{C}$-map (12). Specifically, one can see that the difference of $\rho^{(l)}$ in the adjacent layers is asymptotically of $\rho^{(l)\frac{3}{2}}$, rather than $\rho^{(l)2}$:
Using the above difference equation, we can also derive the nonuniversal metric factor $\kappa$ for the scale-invariant activation functions, whose functional form is visualized in the inset of Fig. 5(a):
$\kappa=\frac{\sqrt{2}(1-a)^{2}}{3(1+a^{2})\pi}.$ (72)
The analysis above potentially provides theoretical foundations of some empirical insights in the literature. First, we can correctly expect that a small leak of $a=0.01$, commonly referred to as leaky ReLU (LReLU) in the literature, is unlikely to have a significant impact on the network performance [80, 81], since the metric factor $\kappa$ changes only by 2%. With larger $a$, however, $\kappa$ noticeably decreases (for instance, it becomes half of the original ReLU at $a=2-\sqrt{3}\sim 0.27$) and the optimal depth for training increases in a reciprocal manner (see the final paragraph of Section IV.2). Consequently, it is possible that the networks with suitably chosen leak works better for a fixed task and other network structures, in particular when the original ReLU network tends to overfit the training data; the superior performance of very leaky ($a\sim 0.18$) ReLU reported by Xu et al. [82] may be seen as a remarkable manifestation of such phenomenology, although the differences in the network architecture must be taken into account for a direct comparison.

### v3

@@ IV.3 Scaling results for the finite networks and different architectures
Figure 3: Finite-size scaling of the order-to-chaos transition in the multilayer perceptrons (1). (a) The order parameter $\rho^{(l)}$ (see Eq. (42)) at a critical point ($(\sigma_{w;c},\sigma_{b})\sim(1.39558,0.3)$ with $\tanh$ activation; $\kappa\sim 0.233498$) for various widths $n$ (ranging from 50 (purple) to 400 (orange)), empirically averaged over $10^{4}$ independent runs. Two orthogonal inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ (with $\|\boldsymbol{x}_{1}\|_{2}=\|\boldsymbol{x}_{2}\|_{2}=1$) of size $n_{\mathrm{in}}=10$ were given. The inset shows the same data rescaled according to the univesal scaling ansatz (43). The black dashed curve indicates the solution (45) of the phenomenological description (44) with $(\kappa,\mu)\sim(0.233498,0.6601)$, where $\mu$ was chosen by fitting the solution to the empirical result for $n=400$. (b) The nonuniversal metric factor $\mu$ as a function of $\kappa$ in the case of $\tanh$ activation, where $\mu$ for each $\kappa$ was estimated from the same fitting as the inset of (a) using the empirical $\rho^{(l)}$ for $n=200$.
The empirical finding above can be heuristically understood by considering a finite-width correction to the mean-field theory. In the case of the finite-width networks, the fourth-order (and other even-order) cumulants come into play, while the third-order (and other odd-order) ones vanish just as the odd-order moments do  [^3] This peculiarity explains why the finite size scaling in the multilayer perceptrons is different from that in the contact process [88] on a complete graph, where one finds the same $\beta$ and $\nu_{\|}$ (Eq. (21)) but the exponent for finite-size scaling (43) is replaced with $-1/2$. If the third-order cumulant remained non-zero, the leading order for the correction would be an order of $n^{-\frac{1}{2}}$.. In the spirit of the asymptotic expansion of the probability distribution [44], this observation indicates that the leading correction to the Gaussian process is an order of $n^{-1}$, the reciprocal of the width. Thus, together with an obvious fact that $\rho=0$ is an absorbing state also for the finite networks, we are led to the following modified phenomenological description:
@@
While it is nowadays well established that the depth-to-width ratio $L/n$ is a key quantity for describing the multilayer perceptrons with finite $n$ [23, 45, 46], the metric factor $\mu$ enriches this insight by providing the means to quantitatively characterize the sensitivity of the network to the width. Specifically, the width $n$ of the network should satisfy $n\gtrsim\mu L$ so that the signal propagation dynamics therein is reasonably well approximated by the infinitely wide limit. This introduces another design consideration for the neural networks. Recalling the relevance of $\kappa L$ for the training dynamics of the networks, one would be tempted to use a critical network with larger $\kappa$ to obtain good generalization with smaller $L$. However, Fig. 3(b) empirically suggests that larger $\kappa$ comes with a cost of larger $\mu$, which imposes an extra computational burden for larger $n$. Hence, if one chooses to operate the network near the infinitely wide limit, one needs to make a trade-off between these two factors for cost-effective training. Theoretically, a more precise formulation of this idea might be achieved by studying the finite-width corrections [47] to the training dynamics, which is beyond the scope of the present work.
@@
although the order parameter can saturate to a non-zero value slightly below the critical point (causing the tilts in the rescaled plot) mainly because the perfect order ($\rho^{(l)}=0$) is virtually unachievable; the imperfection acts as a small fluctuation, to which the setups near the critical point is particularly sensitive  [^4] Other possible (albeit relatively minor) reasons for the deviation from the scaling collapse include the finite-size effect and deviation of the noise from the asymptotic behavior (which is particularly true if the number $c$ of channels is small).. We also checked that the essentially same scenario holds true for the two-dimensional convolutional networks (Fig. 4(b)), where the critical exponents $\beta,\nu_{\|}$ are replaced [56, 57] with
@@
Remarkably, the deviation from the scaling collapse is less prominent in the two-dimensional networks than in their one-dimensional counterparts, which is consistent with our interpretation of the tilts since fluctuations generally become less relevant as the spatial dimensionality goes up [10].
The correspondence to the DP universality class suggest that the signal propagation dynamics in the convolutional networks is highly nontrivial, especially given the notorious difficulty of exactly solving DP [58]. Yet, thanks to the universality of the scaling laws of absorbing phase transitions, semi-quantitative predictions can be gained via simple phenomenological considerations. In particular, the most informative combination of the width $n$ and the depth $L$ for describing the behavior of the critically initialized deep convolutional networks may be changed from $L/n$ to $L/n^{\nu_{\|}/\nu_{\perp}}$ using the corresponding critical exponents
@@ V Discussion
Let us emphasize that successful deep learning can only be achieved via a complicated interplay among various setups, even in one of the simplest cases where NTK describes the training dynamics reasonably well. In addition to the relevance of initialization at criticality [18, 21] and of proper scaling of a learning rate with respect to depth [19], the present work demonstrates the necessity of a more specific, depth-dependent choice of hyperparameters. Furthermore, in the case of a large but finite width, one should also strike a balance between width and depth for efficiency. The fact that all these setups need to be considered simultaneously highlights the major challenge in deep learning, which necessitates extensive study on hyperparameter optimization [61]. To put it the other way around, considerable theoretical insight into the mechanism behind the recent success of deep learning may be obtained by studying the neural networks near the realm of NTK, contrary to a common belief [62, 63].
@@
We foresee some interesting directions for future work. Apart from the ones mentioned in the previous Section  [^5] Namely, (a) derivation of more precise design principles of the neural networks from the universal scaling laws, possibly with prior knowledge about a dataset at hand (Sect. IV.2) and (b) uncovering the details of the trade-off between the width and the depth (Sect. IV.3)., one of the most natural directions is to extend the present framework to more modern architectures [72, 73, 25]. In particular, the skip connections employed in the residual neural networks (ResNet [74]) may be seen as a stimulus kicking the system out of an absorbing state: even if the signals are collapsed during the propagation within a residual block, the skip connection breaks the order before entering the new block. Given that a combination of external drive and dissipation into absorbing states has been conjectured to be a key ingredient of self-organized criticality in physical systems [3, 7], the skip connections may have their unique benefits in deep learning as well. Another, albeit less straightforward, direction is to provide thermodynamic foundations of the speed-accuracy trade-off in deep learning. The deep neural networks in the ordered phase sacrifice speed for accuracy, and vice versa in the chaotic phase. Since the speed-accuracy trade-off has been extensively studied in the context of living systems [75, 76, 77, 78], thermodynamical insights developed therein are likely to be helpful (and indeed, very recently, the trade-off in the diffusion models has been studied from a thermodynamic viewpoint [79]), although pursuing this direction would call for an improved understanding of the thermodynamics of absorbing phase transitions [80, 81]. We believe that further investigation into a parallel between intelligence in living systems and that in artificial neural networks will lead us to a lot of exciting developments, beneficial for both physics and machine learning communities.
@@ Appendix B Scale-invariant activation functions
Figure 5: Universal scaling in the infinitely-wide multilayer perceptrons (1) with scale-invariant activation functions (70). (a) The phase diagram of the signal propagation for ReLU activation ($a=0$). A dashed line is used to indicate the phase boundary because the preactivation variance diverges at the boundary unless $\sigma_{b}=0$ (a black point). (b) The order parameter $\rho^{(l)}$ at the edge of chaos (71) for various combinations of cosine distances $\rho^{(0)}$ and leak parameters $a$ (the resulting $\rho^{(0)}\kappa^{2}$ ranges from $3.6\times 10^{-8}$ (purple) to $6.3\times 10^{-4}$ (orange)), calculated from the mean-field theory (6), (7) and then rescaled according to the scaling ansatz (72). The dashed lines are guides-to-eye for the asymptotic behavior (28) of the scaling function $g$. The inset shows the nonuniversal metric factor $\kappa$ as a function of $a$. (c) The main panel shows the order parameter $\rho^{(l)}$ (see Eq. (42)) at a critical point ($(\sigma_{w;c},\sigma_{b})\sim(\sqrt{2},0)$ with ReLU activation; $\kappa=\sqrt{2}/(3\pi)\sim 0.150053$) for various widths $n$ (ranging from 50 (purple) to 400 (orange)), empirically averaged over $4\times 10^{4}$ independent runs and rescaled according to the universal scaling ansatz (76). Two orthogonal inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ (with $\|\boldsymbol{x}_{1}\|_{2}=\|\boldsymbol{x}_{2}\|_{2}=1$) of size $n_{\mathrm{in}}=10$ were given. The black dashed curve indicates the solution (75) of the phenomenological description (74) with $(\kappa,\mu)\sim(0.150053,0.2676)$, where $\mu$ was chosen by fitting the solution to the empirical result for $n=200$. The inset shows the nonuniversal metric factor $\mu$ as a function of $\kappa$ in the case of scale-invariant activation functions, where $\mu$ for each $\kappa$ was estimated from the same fitting as the main panel using the empirical $\rho^{(l)}$ for $n=200$. (d) The NTK $\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})$ (see Eq. (32)) for various network depths, rescaled according to the universal scaling ansatz (77). The same combinations of $(\rho^{(0)},a)$ as (b) were used. The two horizontal lines are guides-to-eye for the asymptotic behavior as $L\rightarrow\infty$ [21]: the upper one for $\boldsymbol{x}_{1}=\boldsymbol{x}_{2}$ ($\rho^{(0)}=0$), the lower one for $\boldsymbol{x}_{1}\neq\boldsymbol{x}_{2}$ ($\rho^{(0)}>0$).
@@
$(\sigma_{w},\sigma_{b})=\left(\sqrt{\frac{2}{1+a^{2}}},0\right),$ (71)
while it diverges linearly if $\sigma_{b}\neq 0$ (Fig. 5(a)). Note that a special case of Eq. (71) for $a=0$ is nothing but the well-known He initialization [82] for ReLU activation. With this initialization scheme, we have a well-defined iterative $\mathcal{C}$-map (12) and hence we can study the order parameter $\rho^{(l)}$ defined in Eq. (20); Cho and Saul [83] provide technical details on how to analytically deal with the integration appearing in Eq. (12) in the case of ReLU activation. Notice also that $q^{*}$ vanishes for $\sigma_{w}<\sigma_{w;c}$ if $\sigma_{b}=0$. In other words, the neurons within sufficiently deep hidden layers die out [84], which is reminiscent of an absorbing phase transition discussed in the main text.
A natural question is whether one can apply the universal scaling of absorbing phase transitions in the present case, which we will address below. As visualized in Fig. 5(b), we find that the order parameter $\rho^{(l)}$ at the edge of chaos (71) follows the universal scaling ansatz
$\rho^{(l)}\simeq(\kappa l)^{-2}g(\rho^{(0)}(\kappa l)^{2}),\quad\kappa=\frac{\sqrt{2}(1-a)^{2}}{3(1+a^{2})\pi}$ (72)
with a suitable scaling function $g$ having the same asymptotic behavior as Eq. (28), where we dropped the metric factor $\omega$ associated with an initial condition because $\omega=1$ in this case. The difference in the scaling exponent compared to the result (27) in the main text stems from the second-dominant term in the iterative $\mathcal{C}$-map (12). Specifically, one can see that the difference of $\rho^{(l)}$ in the adjacent layers is asymptotically of $\rho^{(l)\frac{3}{2}}$, rather than $\rho^{(l)2}$:
Since the argument for a finite-width correction (see just above Eq. (44)) is not sensitive to a selection of the activation function, the correction can be made in the same manner as the $K^{*}=0$ activation functions
$\frac{\mathrm{d}\rho}{\mathrm{d}l}=-\frac{\mu}{n}\rho-2\kappa\rho^{\frac{3}{2}},$ (74)
although the resulting $l$-dependence of the order parameter $\rho^{(l)}$ and the exponents for the finite-size scaling are different:
$n^{2}\rho^{(l)}\simeq\frac{\mu^{2}\rho_{0}}{\left(2\kappa\sqrt{\rho_{0}}(e^{\frac{\mu l}{2n}}-1)+(\mu/n)e^{\frac{\mu l}{2n}}\right)^{2}},$ (75)
$\rho^{(l)}\simeq n^{-2}f(n^{-1}l).$ (76)
Numerical results shown in Fig. 5(c) validate the finite-size scaling ansatz (76) and suggest that the positive correlation between the two metric factors $\kappa,\mu$ holds true for scale-invariant activation functions (see the inset). Similarly, the universal scaling for NTK holds for small $\rho^{(0)}$ or large $L$ (Fig. 5(d)):
@@
Thus, essentially the same scenario as the $K^{*}=0$ activation functions is applicable to the scale-invariant functions, with a suitable change of the scaling exponents.
The analysis above potentially provides theoretical foundations of some empirical insights in the literature. First, we can correctly anticipate that a small leak of $a=0.01$, commonly referred to as leaky ReLU (LReLU) in the literature, is unlikely to have a significant impact on the network performance [85, 86], since the metric factor $\kappa$ changes only by 2%. With larger $a$, however, $\kappa$ noticeably decreases (for instance, it becomes half of the original ReLU at $a=2-\sqrt{3}\sim 0.27$) and the optimal depth for training increases in a reciprocal manner (see the final paragraph of Sect. IV.2). Consequently, it is possible that the networks with suitably chosen leak works better for a fixed task and other network structures, in particular when the original ReLU network tends to overfit the training data; the superior performance of very leaky ($a\sim 0.18$) ReLU reported by Xu et al. [87] may be seen as a remarkable manifestation of such phenomenology, although the differences in the network architecture must be taken into account for a direct comparison.